# Pseudoentanglement Ain’t Cheap

### v1

@@ 1 Introduction
Recent work [ABF+24] introduced the notion of pseudoentangled quantum states, in analogy with pseudorandomness in classical computation. An ensemble of quantum states is said to be pseudoentangled if states in the ensemble have low entanglement across every bipartition, but they are difficult to distinguish from states with much larger entanglement. A more formal definition is the following:
@@
[ABF+24] showed that pseudoentangled states can be instantiated in polynomial time and logarithmic depth, assuming the existence of quantum-secure one-way functions. So, under a standard cryptographic assumption, there exists an efficient construction of pseudoentanglement. Nevertheless, in applications, we sometimes need constructions that are even simpler and more efficient, due to constraints beyond total gate complexity. For example, [ABF+24] suggested that a construction of “holographic” pseudoentangled states might imply that the AdS/CFT dictionary is hard to compute. However, it remains open to build such pseudoentangled states that are compatible with the laws of AdS/CFT. There are also various other measures of quantum state complexity, beyond circuit depth. For example, a counting argument shows that stabilizer states and free-fermionic states can require super-logarithmic depth, but these states are also “easy” in the sense that their evolutions are efficiently classically simulable [AG04, Val02] and they are efficiently learnable [AG08, Mon17, AG23].
In this work, we study the relationship between pseudoentanglement and non-Clifford complexity. The Clifford group is a remarkably useful object in quantum information that consists of all quantum circuits generated by Hadamard, Phase, and CNOT gates. Clifford gates are almost universal for quantum computing: the addition of any single-qubit non-Clifford gate gives rise to a universal gate set, as shown by Shi [Shi02]. Generally speaking, Clifford gates are “cheaper” than non-Clifford gates, in a sense that can be formalized in a variety of applications. Examples where the cost of a quantum operation is dominated by non-Clifford gates include quantum fault tolerance based on magic state distillation [BK05], near-Clifford classical simulation algorithms [AG04, BG16, RLCK19, BBC+19], and quantum learning algorithms based on the stabilizer formalism [LC22, GIKL23a, GIKL23b, GIKL23c, GIKL23d, LOLH22, LOH23, HG23, CLL23].
@@
As it happens, several of the known pseudoentangled state ensembles are also pseudorandom states ensembles, including the only known instantiations of pseudoentanglement that achieve an optimal gap of $\Theta(n)$ vs. $\omega(\log n)$ [ABF+24, GTB23]. However, [ABF+24] further observed that a pseudoentangled ensemble need not be pseudorandom, nor vice-versa. Hence, it is not clear that the computational resources needed to construct pseudoentangled states mirror those for pseudorandom states.
On the other hand, existing work has made clear that that some amount of non-Cliffordness is needed to generate pseudoentangled states. Fattal, Cubitt, Yamamoto, Bravyi, and Chuang [FCY+04] gave an efficient algorithm for computing the entanglement entropy across any bipartition of a stabilizer state (i.e., a state preparable using Clifford gates only). Combined with Montanaro’s algorithm for learning an unknown stabilizer state [Mon17], this implies that any ensemble {|$\{\ket{\Psi_{k}},\ket{\Phi_{k}}\}_{k}$ of stabilizer states cannot be pseudoentangled.
Our main result is an algorithm with a much stronger guarantee than the combination of [FCY+04, Mon17]. In short, our algorithm estimates the entanglement entropy of an unknown quantum state across any cut of qubits, where the accuracy of the estimate scales with the number of Pauli operators that stabilize the state (i.e., the number of Pauli operators for which the state is a $+1$-eigenvector).
@@
We remark that the set of states for which this is applicable is rather large and includes states prepared by a Clifford circuit with up to $\lfloor\frac{t}{2}\rfloor$ auxiliary single-qubit non-Clifford gates (such as the $T$-gate). When the algorithm is run on a stabilizer state (i.e., when $t=0$), then our algorithm outputs the entanglement entropy across any cut of qubits exactly, recovering the aforementioned result of [FCY+04, Mon17]. On the other hand, the bounds degrade as $t$ becomes too close to $n$.
@@
While finalizing this work, we became aware of independent and concurrent work by Gu, Oliveiro, and Leone [GOL24]. Their [GOL24, Lemma 7] resembles our Theorem 3.1. While their result applies to all $\alpha$-Rényi entanglement entropies, their lower bound is looser than ours for the case of von Neumann entropy (i.e., $\alpha=1$). As a result, our Algorithm 1 produces an estimate of von Neumann entropy that is a factor of $2$ tighter than the output of their [GOL24, Algorithm 3]. In addition to algorithms for estimating $\alpha$-Rényi entanglement entropies, [GOL24] proves other results that probe the relationship between magic and entanglement.
@@ 3 Entanglement Entropy Bounds
We prove upper and lower bounds on the entanglement entropy across any cut of qubits for any $n$-qubit quantum state |$\ket{\psi}$. The quality of our bounds depends on $\mathrm{Weyl}(\ket{\psi})$. For example, if $\dim(\mathrm{Weyl}(\ket{\psi}))=0$, our bounds become trivial, and, if $\dim(\mathrm{Weyl}(\ket{\psi}))=n$ (i.e., |$\ket{\psi}$ is a stabilizer state), our bounds are tight, recovering the main result of [FCY+04].
@@ 3.2 Proof of Lower Bound
Our last step is to prove that $C^{B}(C^{A}(T))_{B^{\prime}}=\{0\}$, which will imply that the entanglement across $(A^{\prime},B^{\prime})$ is $\dim(T)/2$ by [FCY+04, Eq. 1]. First recall that $S\cap\left(\mathbb{F}_{2}^{2n}\right)_{B}=\{0\}$ by construction, which implies $T\cap\left(\mathbb{F}_{2}^{2n}\right)_{B}=\{0\}$ because $T\subseteq S$. Next, we note that $C^{A}$ has no effect on $\left(\mathbb{F}_{2}^{2n}\right)_{B}$ since its action is local to $A$. Furthermore, $C^{B}$ simply permutes $\left(\mathbb{F}_{2}^{2n}\right)_{B}$ (and cannot map elements into $(\mathbb{F}_{2}^{2n})_{B})$. Hence,
@@ 4 The Algorithm
1
2Perform Bell difference sampling to draw $\frac{2\ln(1/\delta)+4n}{\varepsilon^{2}}$ samples from $q_{\psi}$.
3Let $S$ be the symplectic complement of the subspace spanned by the samples.
4Let $r\coloneqq\begin{cases}0&\dim(S)=n-k,\\ \varepsilon n+\mathsf{H}(\varepsilon)&\dim(S)>n-k.\end{cases}$
5Let $u\coloneqq\min\{|A|-\dim S_{A},|B|-\dim S_{B}\}+r$.
6Let $\ell\coloneqq\max\{\dim S-\dim S_{B}-|A|,\dim S-\dim S_{A}-|B|\}-r$.
7return $(\ell,u)$

### v2

@@ 1 Introduction
Recent work [${}^{+}$24] introduced the notion of pseudoentangled quantum states, in analogy with pseudorandomness in classical computation. An ensemble of quantum states is said to be pseudoentangled if states in the ensemble have low entanglement across every bipartition, but they are difficult to distinguish from states with much larger entanglement. A more formal definition is the following:
@@
[${}^{+}$24] showed that pseudoentangled states can be instantiated in polynomial time and logarithmic depth, assuming the existence of quantum-secure one-way functions. So, under a standard cryptographic assumption, there exists an efficient construction of pseudoentanglement. Nevertheless, in applications, we sometimes need constructions that are even simpler and more efficient, due to constraints beyond total gate complexity. For example, [${}^{+}$24] suggested that a construction of “holographic” pseudoentangled states might imply that the AdS/CFT dictionary is hard to compute. However, it remains open to build such pseudoentangled states that are compatible with the laws of AdS/CFT. There are also various other measures of quantum state complexity, beyond circuit depth. For example, a counting argument shows that stabilizer states and free-fermionic states can require super-logarithmic depth, but these states are also “easy” in the sense that their evolutions are efficiently classically simulable [AG04, Val02] and they are efficiently learnable [AG08, Mon17, AG23].
In this work, we study the relationship between pseudoentanglement and non-Clifford complexity. The Clifford group is a remarkably useful object in quantum information that consists of all quantum circuits generated by Hadamard, Phase, and CNOT gates. Clifford gates are almost universal for quantum computing: the addition of any single-qubit non-Clifford gate gives rise to a universal gate set, as shown by Shi [Shi02]. Generally speaking, Clifford gates are “cheaper” than non-Clifford gates, in a sense that can be formalized in a variety of applications. Examples where the cost of a quantum operation is dominated by non-Clifford gates include quantum fault tolerance based on magic state distillation [BK05], near-Clifford classical simulation algorithms [AG04, BG16, RLCK19, ${}^{+}$19], and quantum learning algorithms based on the stabilizer formalism [LC22, GIKL23a, GIKL23b, GIKL23c, GIKL23d, LOLH22, LOH23, HG23, CLL23].
@@
As it happens, several of the known pseudoentangled state ensembles are also pseudorandom states ensembles, including the only known instantiations of pseudoentanglement that achieve an optimal gap of $\Theta(n)$ vs. $\omega(\log n)$ [${}^{+}$24, GTB23]. However, [${}^{+}$24] further observed that a pseudoentangled ensemble need not be pseudorandom, nor vice-versa. Hence, it is not clear that the computational resources needed to construct pseudoentangled states mirror those for pseudorandom states.
On the other hand, existing work has made clear that that some amount of non-Cliffordness is needed to generate pseudoentangled states. Fattal, Cubitt, Yamamoto, Bravyi, and Chuang [${}^{+}$04] gave an efficient algorithm for computing the entanglement entropy across any bipartition of a stabilizer state (i.e., a state preparable using Clifford gates only). Combined with Montanaro’s algorithm for learning an unknown stabilizer state [Mon17], this implies that any ensemble {|$\{\ket{\Psi_{k}},\ket{\Phi_{k}}\}_{k}$ of stabilizer states cannot be pseudoentangled.
Our main result is an algorithm with a much stronger guarantee than the combination of [${}^{+}$04, Mon17]. In short, our algorithm estimates the entanglement entropy of an unknown quantum state across any cut of qubits, where the accuracy of the estimate scales with the number of Pauli operators that stabilize the state (i.e., the number of Pauli operators for which the state is a $+1$-eigenvector).
@@
We remark that the set of states for which this is applicable is rather large and includes states prepared by a Clifford circuit with up to $\lfloor\frac{t}{2}\rfloor$ auxiliary single-qubit non-Clifford gates (such as the $T$-gate). When the algorithm is run on a stabilizer state (i.e., when $t=0$), then our algorithm outputs the entanglement entropy across any cut of qubits exactly, recovering the aforementioned result of [${}^{+}$04, Mon17]. On the other hand, the bounds degrade as $t$ becomes too close to $n$.
@@
While finalizing this work, we became aware of independent and concurrent work by Gu, Oliveiro, and Leone [GOL24]. Their [GOL24, Lemma 7] resembles our Theorem 3.1. However, we only prove bounds for the von Neumann entropy, whereas [GOL24] prove similar bounds for any $\alpha$-Rényi entanglement entropy, which captures the von Neumann entropy as a special case.
@@ 3 Entanglement Entropy Bounds
We prove upper and lower bounds on the entanglement entropy across any cut of qubits for any $n$-qubit quantum state |$\ket{\psi}$. The quality of our bounds depends on $\mathrm{Weyl}(\ket{\psi})$. For example, if $\dim(\mathrm{Weyl}(\ket{\psi}))=0$, our bounds become trivial, and, if $\dim(\mathrm{Weyl}(\ket{\psi}))=n$ (i.e., |$\ket{\psi}$ is a stabilizer state), our bounds are tight, recovering the main result of [${}^{+}$04].
@@ 3.2 Proof of Lower Bound
Our last step is to prove that $C^{B}(C^{A}(T))_{B^{\prime}}=\{0\}$, which will imply that the entanglement across $(A^{\prime},B^{\prime})$ is $\dim(T)/2$ by [${}^{+}$04, Eq. 1]. First recall that $S\cap\left(\mathbb{F}_{2}^{2n}\right)_{B}=\{0\}$ by construction, which implies $T\cap\left(\mathbb{F}_{2}^{2n}\right)_{B}=\{0\}$ because $T\subseteq S$. Next, we note that $C^{A}$ has no effect on $\left(\mathbb{F}_{2}^{2n}\right)_{B}$ since its action is local to $A$. Furthermore, $C^{B}$ simply permutes $\left(\mathbb{F}_{2}^{2n}\right)_{B}$ (and cannot map elements into $(\mathbb{F}_{2}^{2n})_{B})$. Hence,
@@ 4 The Algorithm
Perform Bell difference sampling to draw $\frac{2\ln(1/\delta)+4n}{\varepsilon^{2}}$ samples from $q_{\psi}$. Let $S$ be the symplectic complement of the subspace spanned by the samples. Let $r\coloneqq\begin{cases}0&\dim(S)=n-k,\\ \varepsilon n+\mathsf{H}(\varepsilon)&\dim(S)>n-k.\end{cases}$ Let $u\coloneqq\min\{|A|-\dim S_{A},|B|-\dim S_{B}\}+r$. Let $\ell\coloneqq\max\{\dim S-\dim S_{B}-|A|,\dim S-\dim S_{A}-|B|\}-r$. return $(\ell,u)$